# Are Neuromorphic Architectures Inherently Privacy-preserving? An Exploratory Study

### v1

@@ Abstract.
While machine learning (ML) models are becoming mainstream, including in critical application domains, concerns have been raised about the increasing risk of sensitive data leakage. Various privacy attacks, such as membership inference attacks (MIAs), have been developed to extract data from trained ML models, posing significant risks to data confidentiality. While the mainstream work in the ML community considers traditional Artificial Neural Networks (ANNs) as the default neural model, neuromorphic architectures, such as Spiking Neural Networks (SNNs) have recently emerged as an attractive alternative mainly due to their significantly low power consumption. These architectures process information through discrete events, i.e., spikes, to mimic the functioning of biological neurons in the brain. While the privacy issues have been extensively investigated in the context of traditional ANNs, they remain largely unexplored in neuromorphic architectures, and little work has been dedicated to investigate their privacy-preserving properties.
In this paper, we investigate the question whether SNNs have inherent privacy-preserving advantage. Specifically, we investigate SNNs’ privacy properties through the lens of MIAs across diverse datasets, comparatively with ANNs . We explore the impact of different learning algorithms (surrogate gradient and evolutionary learning), programming frameworks (snnTorch, TENNLab, and LAVA), and various parameters on the resilience of SNNs against MIA. Our experiments reveal that SNNs demonstrate consistently superior privacy preservation compared to ANNs, with evolutionary algorithms further enhancing their resilience. For example, on the CIFAR-10 dataset, SNNs achieve an AUC as low as 0.59 compared to 0.82 for ANNs, and on CIFAR-100, SNNs maintain a low AUC of 0.58, whereas ANNs reach 0.88. Furthermore, we investigate the privacy-utility trade off through Differentially Private Stochastic Gradient Descent (DPSGD) observing that SNNs incur a notably lower accuracy drop than ANNs under equivalent privacy constraints.
@@ 1. Introduction
As ML systems become more sophisticated and widespread, individuals are increasingly relying on these systems, entrusting them with personal and professional data. Consequently, the risk of sensitive information exposure is growing significantly in multiple sectors (Bertino, 2016) including healthcare (Abouelmehdi et al., 2017), finance (Tripathi and Mukhopadhyay, 2020), national security (Thuraisingham, 2002), education (Florea and Florea, 2020) and consumer services (Lee et al., 2015). It is particularly alarming in fields such as healthcare, where the confidentiality of patient data is extremely sensitive as a breach could result in severe personal and financial implications, affecting patient care and institutional credibility (Luo et al., 2018; JM et al., 2018; U.S. Department of Health & Human Services, 2024). In finance, the integrity of financial transactions and records is fundamental for maintaining market stability and preventing fraud (Yu and He, 2021), while in national security , safeguarding classified information is essential to protect national interests and prevent threats to public safety (Kim et al., 2010).
This has led to the development of various privacy attacks targeting ML models to extract sensitive information, including Model Inversion Attacks (Fredrikson et al., 2015), Attribute Inference Attacks (Gong and Liu, 2018), Model Stealing Attacks (Juuti et al., 2019), and Membership Inference Attacks (MIAs) (Shejwalkar et al., 2021). Among these, MIAs are particularly prominent as a significant threat to data privacy (Salomon, 2012), wherein an adversary seeks to ascertain if a specific data point was part of the dataset used to train the model. This intrusion risks exposing sensitive information about individuals in the training dataset, potentially compromising personal data confidentiality (Shokri et al., 2017).
While most research on privacy attacks has concentrated on ANNs, there is limited exploration within neuromorphic computing, particularly with Spiking Neural Networks (SNNs). Designed to emulate the dynamic behavior of biological neurons (Ghosh-Dastidar and Adeli, 2009), SNNs process information through discrete, temporally encoded spikes (Roy et al., 2019), enabling them to handle time-sensitive data efficiently (Hong et al., 2017). Their suitability for edge computing (Shi et al., 2016) and resource-constrained environments further enhances their value, as SNNs effectively process real-world spatiotemporal patterns (Wang et al., 2018). This capability positions SNNs as a promising alternative to traditional neural networks for applications requiring dynamic, real-time data processing.
@@
Investigating the resilience of SNNs against MIAs, our experimental results consistently demonstrated that SNNs exhibit higher resilience to MIAs across the datasets including MNIST, F-MNIST, Iris, Breast Cancer, CIFAR-10, CIFAR-100, and ImageNet. This is evidenced by the lower Area Under the Curve (AUC) values for the Receiver Operating Characteristic (ROC) curves in SNNs compared to their ANN counterparts. Furthermore, our exploration domain encompassed various learning algorithms (surrogate gradient-based and evolutionary learning), programming frameworks (snnTorch, TENNLab, and LAVA), and a wide range of parameters within them, providing a comprehensive analysis of the factors influencing the inherent privacy-preserving properties of SNNs. This in-depth exploration revealed that evolutionary learning algorithms have shown to boost this resilience more effectively compared to the gradient-based methods.
@@
This paper offers the following notable findings in the field of data privacy, particularly in the context of SNNs:
@@
  • SNNs exhibit higher resilience against MIAs compared to ANNs, with lower AUC scores on CIFAR-10 (SNN: 0.59 vs. ANN: 0.82) and CIFAR-100 (SNN: 0.58 vs. ANN: 0.88), highlighting their potential as a more secure alternative in privacy-sensitive applications.
@@
  • Evolutionary learning algorithms outperform gradient-based methods in MIA resilience, maintaining a consistent AUC of  0.50 across all parameters for Iris and Breast Cancer datasets, compared to 0.57 and 0.55 AUC scores for gradient-based algorithms, respectively.
@@
  • Privacy-utility tradeoff analysis revealing that SNNs incur a lower accuracy drop compared to ANNs when applying DPSGD: For F-MNIST, with privacy guarantees ranging from 0.22 to 2.00, the average accuracy drop is 12.87% for SNNs, significantly lower than the 19.55% drop observed in ANNs.
@@
It should be emphasized that while this investigation highlights SNNs’ enhanced privacy characteristics, these findings are specifically contextualized within privacy-preservation applications. The architectural properties of SNNs that facilitate efficient hardware implementation and reduced computational overhead make them particularly appealing for resource-constrained environments. However, these advantages should not be interpreted as a general superiority of SNNs over ANNs across all applications. Rather, this work is motivated by the intuition that SNNs’ unique information processing mechanisms may offer specific advantages in privacy preservation, warranting systematic investigation in this particular domain.
2. Related Works
As machine learning systems handle increasingly sensitive data, the potential for privacy violations grows. Li et al.  (Liu et al., 2021a) categorize these privacy challenges into two primary areas: privacy attacks and privacy-preserving techniques. Privacy attacks have emerged as a critical concern in ML due to the growing realization that models can act as unintended leak vectors. These attacks can broadly be classified into different types, such as model inversion attacks, model extraction attacks, MIA. Model inversion attacks (Fredrikson et al., 2015) reconstruct input data from model outputs, while model extraction attacks  (Juuti et al., 2019) aim to replicate a model’s functionality without direct access to its architecture or parameters. Among these, MIAs have gained significant attention due to their ability to infer whether a specific data point was used in training a model. According to the survey conducted by  (Hu et al., 2022), MIAs were first proposed in the context of genomic data by Homer et al. (Homer et al., 2008) where an attacker could infer if an individual’s genome was part of a dataset based on summary statistics. Later, Shokri et al.(Shokri et al., 2017) introduced the first systematic MIA framework, showing how adversaries could use shadow models to infer training data membership. Salem et al.  (Salem et al., 2018) reduced the complexity by demonstrating that a single shadow model can perform well compared to using multiple models, and they introduced metric-based attacks that rely on confidence scores and entropy without the need for identical data distribution between shadow and target models. Nasr et al. (Nasr et al., 2019) further expanded MIA into white-box settings, demonstrating that attackers with access to internal model parameters can perform even more effective MIAs. Melis et al. (Melis et al., 2019) extended MIA to federated learning, showing vulnerabilities even in distributed learning settings, where multiple parties collaboratively train a model. Song and Mittal (Song and Mittal, 2021) highlighted the increased privacy risks in generative models such as GANs, where membership inference attacks could be carried out on synthetic data generators. Recent work by Ilyas et al. introduced LiRA (Likelihood Ratio Attack)  (Carlini et al., 2022), a method that further improves the accuracy of MIAs by leveraging confidence scores more effectively to distinguish between training and non-training data points.In 2024, Zarifzadeh et al. (Zarifzadeh et al., 2024) introduced RMIA, a high-power membership inference attack that outperforms prior methods like LiRA and Attack-R, demonstrating superior robustness, particularly at low FPRs, using likelihood ratio tests.
To counteract these privacy attacks, several privacy-preserving techniques have been developed.These techniques range from cryptographic approaches like homomorphic encryption(HE) (Yi et al., 2014) and secure multi-party computation(SMPC) (Zhao et al., 2019) to learning-based defenses such as model obfuscation (Guo et al., 2021) and knowledge distillation (Gou et al., 2021). However, these methods often suffer from computational inefficiencies, particularly in large-scale systems. Another approach is Federated Learning (FL) (Li et al., 2020), which enables collaborative model training without sharing raw data, but remains vulnerable to attacks like MIA. Among these, Differential Privacy (DP) (Dwork, 2006) has gained prominence due to its strong theoretical guarantees and practical applicability in machine learning settings.It provides a robust framework for protecting individual data points by introducing noise during computations. In ML, this concept has been adapted through various algorithms, with DPSG) (Song et al., 2013) being the most prominent which applies DP principles by adding noise to the gradient updates during training, offering a practical way to maintain privacy while training large models without significantly compromising accuracy.
While much of the data privacy research has centered around ANNs, expanding these investigations to SNNs is necessary. SNNs not only offer performance levels comparable to ANNs but also exhibit superior energy efficiency and hardware integration capabilities, positioning them as promising candidates for exploring inherent privacy features. Although privacy attacks on neuromorphic architectures remain underexplored, existing studies have yet to confirm SNNs’ potential resistance to such threats. However, significant strides have been made in privacy-preserving techniques within the neuromorphic domain. For instance, recent efforts by Han et al.  (Han et al., 2023) focus on developing privacy-preserving methods for SNNs, particularly utilizing FL and DP to address both computational efficiency and privacy challenges. Li et al.(Li et al., 2023) introduced a framework that combines Fully Homomorphic Encryption (FHE) with SNNs, enabling encrypted inference while preserving SNNs’ energy efficiency and computational advantages. Similarly, Nikfam et al.(Nikfam et al., 2023) developed an HE framework tailored for SNNs, offering enhanced accuracy over DNNs under encryption, while carefully balancing computational efficiency. Additionally, Safronov et al.(Kim et al., 2022) proposed PrivateSNN, a privacy-preserving framework for SNNs that employs differential privacy to mitigate membership inference attacks, maintaining the energy-efficient nature of SNNs.
@@ 3. Background
Figure 1. Schematic representation of the experimental flow. The framework begins by comparing the privacy risks in ANNs and SNNs through MIA(yellow block). The SNN-specific exploration(green block) space evaluates privacy resilience using surrogate gradient algorithms and evolutionary algorithms with varying parameters. Finally, both ANNs and SNNs undergo DPSGD algorithm(blue block) to assess the privacy-utility tradeoff
@@ 3.1. Neuromorphic Architecture
Neuromorphic architectures  (Marković et al., 2020) are designed to replicate the neural structures and functionalities of the biological brain, offering an alternative to traditional von Neumann computing systems  (Aspray, 1990). These architectures integrate memory and processing units, enabling massive parallel processing  (Nassi and Callaway, 2009) and event-driven computation  (Hwang and Wu, 2000) .By modeling neurons and synapses that communicate via discrete spikes or events, neuromorphic systems can process data asynchronously, reducing power consumption and latency  (Liu et al., 2021b). Hardware implementations like IBM’s TrueNorth  (Akopyan et al., 2015) and Intel’s Loihi (Davies et al., 2018) demonstrate the potential for scalable, energy-efficient architectures capable of performing complex tasks with a structure that closely resembles biological neural networks.
SNNs are a core component of neuromorphic systems offering a distinct approach compared to traditional ANNs. In SNNs, neurons operate by accumulating charge over time, firing only when a threshold is reached, thereby producing discrete spikes. This introduces a temporal dimension to information processing, as the timing of these spikes encodes key information. SNN models range from basic integrate-and-fire neurons  (Auge et al., 2021), where spikes are triggered by charge accumulation, to more complex, biologically realistic models like the Hodgkin-Huxley model, which captures finer details of neuronal dynamics. This spike-based communication not only mirrors biological neurons but also enables asynchronous data flow, making SNNs particularly suitable for neuromorphic hardware that thrives on event-driven processing. The inclusion of temporal dynamics and synaptic weights in SNNs allows for a more adaptable learning process, which is essential for tasks requiring real-time decision-making and low-latency responses  (Schuman et al., 2022a).
The unique characteristics of SNNs, particularly their spiking-based encoding and asynchronous processing, introduce complexities that may impact their vulnerability to privacy attacks. Traditional ANNs are prone to attacks such as Membership Inference Attacks (MIAs), where adversaries can exploit deterministic output patterns and smooth gradients to infer the presence of specific data points in the training set. In contrast, SNNs operate with discrete, non-differentiable spikes  (Zhang et al., 2018), where the timing and stochastic nature of spike generation add variability to the output. Neurons in SNNs accumulate charge from inputs and fire based on probabilistic thresholds, which can lead to diverse spike patterns even for similar inputs. This inherent variability and the discontinuous nature of SNN activations complicate the tracing of input-output relationships, making it harder for adversaries to identify patterns necessary for MIAs. Although this suggests a potential for increased privacy resilience, especially when combined with techniques like DPSGD, this remains a hypothesis. Our investigation explores whether SNNs, due to their spike-based processing, exhibit lower performance degradation under privacy constraints when compared to ANNs.
@@ 3.2. Membership Inference Attack (MIA)
MIA is a privacy attack that enables adversaries to infer the presence of individual data in a machine learning model’s training dataset. This attack exploits the differences in the model’s behavior when queried with data points that were part of the training set versus those that were not  (Rahman et al., 2018). Typically, models tend to exhibit higher confidence or different error patterns for samples they have encountered during training, compared to new unseen data  (Nguyen et al., 2015). By analyzing these patterns, attackers can infer sensitive information about the training data, which can lead to privacy breaches  (De Cristofaro, 2020), especially in applications dealing with confidential or personal data.
MIA involves training attack models or employing statistical tests to distinguish between the model responses on training and non-training data. The success of these attacks largely depends on the model’s overfitting to its training data and the distinctiveness of the model’s responses to individual samples. MIAs not only breach the privacy of data but also expose weaknesses in a model’s ability to generalize (Gomm et al., 2000).
@@ 3.3. Differentially Private Stochastic Gradient Descent (DPSGD)
Differential Privacy (DP)  (Dwork, 2006) is a mathematical framework that aims to ensure that the output of a computation remains nearly unchanged irrespective of whether an individual’s data is included in the dataset, thus offering strong privacy guarantees. It provides a framework to quantify the impact on privacy when a single data point in a dataset is changed. The privacy parameters $\epsilon$ (privacy budget) and $\delta$ (delta parameter) control this impact. The core principle is expressed in the inequality:
@@
where $D$ and $D^{\prime}$ are datasets differing by one element, $M$ is a randomized algorithm, and $S$ is a subset of potential outputs. This inequality limits the probability of $M$ producing an output within $S$ when applied to $D$ is limited by the exponential of $\epsilon$. This probability is then multiplied by the probability of $M$ producing the same output from $D^{\prime}$. Additionally, it is adjusted by a small term $\delta$. Here, $\epsilon$ modulates the sensitivity of the output to input changes—lower values imply stronger privacy by reducing the allowable change in output probabilities. $\delta$, ideally a small value near zero, accounts for the rare cases where this privacy guarantee might fail, thus ensuring that changes to any single data point in $D$ minimally affect the output, reinforcing data privacy across the dataset.
DPSGD utilizes the principles of DP by incorporating a mechanism of noise addition directly into the gradient descent process used for training machine learning models. In DPSGD, the key step involves perturbing the gradients computed during each training iteration with noise that is calibrated to the sensitivity of the function being optimized. This sensitivity measures how much the output of a function can change in response to changes in its input, which in the context of machine learning, translates to how much a single training example can influence the overall model parameters. The noise added is typically drawn from a Gaussian distribution  (Kassam, 2012), scaled $\epsilon$ and the desired level of privacy guarantee, $\delta$.  (OpenAI, 2024; Anthropic, 2024)
The function of DPSGD can be expressed mathematically as:

### v2

@@ Abstract.
While machine learning (ML) models are becoming mainstream, including in critical application domains, concerns have been raised about the increasing risk of sensitive data leakage. Various privacy attacks, such as membership inference attacks (MIAs), have been developed to extract data from trained ML models, posing significant risks to data confidentiality. While the predominant work in the ML community considers traditional Artificial Neural Networks (ANNs) as the default neural model, neuromorphic architectures, such as Spiking Neural Networks (SNNs) have recently emerged as an attractive alternative mainly due to their significantly low power consumption. These architectures process information through discrete events, i.e., spikes, to mimic the functioning of biological neurons in the brain. While the privacy issues have been extensively investigated in the context of traditional ANNs, they remain largely unexplored in neuromorphic architectures, and little work has been dedicated to investigate their privacy preserving properties. In this paper, we investigate the question whether SNNs have inherent privacy preserving advantage. Specifically, we investigate SNNs’ privacy properties through the lens of MIAs across diverse datasets, in comparison with ANNs. We explore the impact of different learning algorithms (surrogate gradient and evolutionary learning), programming frameworks (snnTorch, TENNLab, and LAVA), and various parameters on the resilience of SNNs against MIA. Our experiments reveal that SNNs demonstrate consistently superior privacy preservation compared to ANNs, with evolutionary algorithms further enhancing their resilience. For example, on the CIFAR-10 dataset, SNNs achieve an AUC as low as 0.59 compared to 0.82 for ANNs, and on CIFAR-100, SNNs maintain a low AUC of 0.58, whereas ANNs reach 0.88. Furthermore, we investigate the privacy-utility trade off through Differentially Private Stochastic Gradient Descent (DPSGD) observing that SNNs incur a notably lower accuracy drop than ANNs under equivalent privacy constraints.
Membership Inference Attack, Spiking Neural Network, Differentially Private Stochastic Gradient Descent, Neuromorphic Computing
@@ 1. Introduction
As ML systems become more sophisticated and widespread, individuals are increasingly relying on these systems, entrusting them with personal and professional data. Consequently, the risk of sensitive information exposure (Salomon, 2012) is growing significantly in multiple sectors (Bertino, 2016) including healthcare (Abouelmehdi et al., 2017), finance (Tripathi and Mukhopadhyay, 2020), national security (Thuraisingham, 2002), education (Florea and Florea, 2020) and consumer services (Lee et al., 2015). It is particularly alarming in fields such as healthcare, where the confidentiality of patient data is extremely sensitive, as a breach could result in severe personal and financial implications, which can affect patient care and institutional credibility (Luo et al., 2018; JM et al., 2018; U.S. Department of Health & Human Services, 2024). In finance, the integrity of financial transactions and records is fundamental for maintaining market stability and preventing fraud (Yu and He, 2021), while in national security, safeguarding classified information is essential to protect national interests and prevent threats to public safety (Kim et al., 2010).
This has led to the development of various privacy attacks targeting ML models to extract sensitive information, including Model Inversion Attacks (Fredrikson et al., 2015), Attribute Inference Attacks (Gong and Liu, 2018), Model Stealing Attacks (Juuti et al., 2019), and Membership Inference Attacks (MIAs) (Shejwalkar et al., 2021). In MIAs, an adversary seeks to ascertain if a specific data point was part of the dataset used to train the model. This intrusion risks exposing classified information about individuals in the training dataset, potentially compromising personal data confidentiality (Shokri et al., 2017).
Designed to replicate the dynamic behavior of biological neurons (Ghosh-Dastidar and Adeli, 2009), SNNs process information through discrete, temporally encoded spikes (Roy et al., 2019), enabling them to handle time-sensitive data efficiently (Hong et al., 2017). Their suitability for edge computing (Shi et al., 2016) and resource-constrained environments further enhances their value, as SNNs effectively process real-world spatiotemporal patterns (Wang et al., 2018). This capability positions SNNs as a promising alternative to traditional neural networks for applications requiring dynamic, real-time data processing. While the security of SNNs has been investigated in the literature (El-Allami et al., 2021; Venceslai et al., 2020), relatively little attention has been given to their privacy-preserving capabilities.
@@
Investigating the resilience of SNNs against MIAs, our experimental results consistently demonstrate that SNNs exhibit higher resilience to MIAs across the datasets including MNIST, F-MNIST, Iris, Breast Cancer, CIFAR-10, and CIFAR-100. This is evidenced by the lower Area Under the Curve(AUC) values for the Receiver Operating Characteristic(ROC) curves in SNNs compared to their ANN counterparts. Furthermore, our exploration domain encompasses various learning algorithms (surrogate gradient-based and evolutionary learning), programming frameworks (snnTorch, TENNLab, and LAVA), and a wide range of parameters within them, providing a comprehensive analysis of the factors influencing the inherent privacy preserving properties of SNNs. This in-depth exploration indicated that evolutionary learning algorithms shown to boost this resilience more effectively compared to the gradient based methods.
@@
This paper offers the following notable contributions in the field of data privacy, particularly in the context of SNNs:
@@
  • SNNs exhibit higher resilience against MIAs compared to ANNs, with lower AUC scores on CIFAR-10 (SNN: 0.59 vs. ANN: 0.82) and CIFAR-100 (SNN: 0.58 vs. ANN: 0.88), highlighting their potential as a more secure alternative in privacy sensitive applications.
@@
  • Evolutionary learning algorithms outperform gradient based methods in MIA resilience, maintaining a consistent AUC of  0.50 across all parameters for Iris and Breast Cancer datasets, compared to 0.57 and 0.55 AUC scores for gradient-based algorithms, respectively.
@@
  • Privacy-utility trade off analysis revealing that SNNs incur a lower accuracy drop compared to ANNs when applying DPSGD: For F-MNIST, with privacy guarantees ranging from 0.22 to 2.00, the average accuracy drop is 12.87% for SNNs comparatively lower than the 19.55% drop observed in ANNs.
@@
We emphasize that while this investigation highlights SNNs’ enhanced privacy characteristics, our findings specifically address privacy preservation applications. The architectural properties of SNNs that enable efficient hardware implementation and reduce computational overhead make them particularly appealing for resource constrained environments. However, these findings do not suggest an overall superiority of SNNs over ANNs across different application domains. Rather, we base this work on the intuition that SNNs’ unique information processing mechanisms may offer specific advantages in privacy preservation, which warrants systematic investigation in this particular domain. The implementation of all experiments conducted in this study is publicly available at this repository.
@@ 2. Background
Figure 1. Architectural overview of the experimental framework. The methodology has 3 components: (i) comparative privacy assessment between ANNs and SNNs through MIA (yellow), (ii) exploration of SNN specific privacy characteristics considering surrogate gradient and evolutionary algorithms (green), and (iii) Comparative evaluation of privacy-utility trade offs through DPSGD (blue)
[…]
@@ 2.1. Neuromorphic Architecture
Neuromorphic architectures  (Marković et al., 2020) are designed to mimic the neural structures and functionalities of the biological brain, offering an alternative to the traditional von Neumann computing systems  (Aspray, 1990). These architectures integrate memory and processing units, enabling massive parallel processing  (Nassi and Callaway, 2009) and event driven computation  (Hwang and Wu, 2000). By modeling neurons and synapses that communicate via discrete spikes or events  (Liu et al., 2021b), neuromorphic systems can process data asynchronously, reducing power consumption and latency. Hardware implementations like IBM’s TrueNorth  (Akopyan et al., 2015) and Intel’s Loihi  (Davies et al., 2018) demonstrate the potential for scalable, energy efficient architectures capable of performing complex tasks with a structure resembling biological neural networks.
SNNs constitute the foundational architecture of neuromorphic systems, operating through a mechanism where neurons accumulate membrane potential over time and generate discrete spike events upon reaching a threshold potential. This temporal encoding paradigm fundamentally differentiates SNNs from traditional neural networks, as information is encoded in both spike timing and frequency. The computational models of SNNs span from elementary integrate-and-fire neurons  (Auge et al., 2021) to sophisticated biologically inspired implementations such as the Hodgkin Huxley model, which incorporates detailed ionic conductance dynamics. The event driven nature of spike-based computation intrinsically aligned with neuromorphic hardware architectures, enabling efficient asynchronous processing. This integration of temporal dynamics and synaptic plasticity in SNNs facilitates adaptive learning mechanisms particularly suited for applications demanding real time processing and minimal latency  (Schuman et al., 2022a).
The architectural distinctiveness of SNNs, characterized by spike based information encoding and asynchronous processing, introduces computational complexities that potentially influence their susceptibility to privacy attacks. While traditional ANNs demonstrate vulnerability to Membership Inference Attacks through their deterministic output patterns and continuous gradients, SNNs operate via discrete, non-differentiable spike events  (Zhang et al., 2018). The temporal dynamics of spike generation and the stochastic nature of neuronal activation contribute to output variability. The membrane potential accumulation in SNN neurons, coupled with probabilistic firing thresholds, generates diverse spike patterns for comparable inputs. This intrinsic variability and the discontinuous activation characteristics inherent to SNNs potentially obscure input-output relationships, thereby complicating the pattern recognition essential for successful MIAs. Although these properties, particularly when combined with differential privacy techniques like DPSGD, suggest enhanced privacy preservation capabilities, such assumptions require rigorous validation. This investigation examines whether the spike based computational paradigm of SNNs exhibits superior resilience to performance degradation under privacy constraints compared to traditional ANNs.
@@ 2.2. Membership Inference Attack (MIA)
constitute a class of privacy vulnerabilities that enable adversaries to determine whether specific data points were used in a model’s training dataset. These attacks exploit differential behavioral patterns in model responses between training and non-training samples  (Rahman et al., 2018). Neural networks typically demonstrate heightened prediction confidence and distinctive error distributions for previously encountered training samples compared to unseen data  (Nguyen et al., 2015). Through systematic analysis of these response characteristics, adversaries can extract sensitive training set information, potentially compromising data privacy  (De Cristofaro, 2020) in domains handling confidential data. The implementation of MIAs encompasses both the development of specialized attack models and the application of statistical inference methods to differentiate between model responses to training and non-training samples. The efficacy of these attacks correlates strongly with the degree of model overfitting and the distinctiveness of individual sample responses. Beyond immediate privacy implications, MIAs serve as indicators of model generalization deficiencies  (Gomm et al., 2000), highlighting vulnerabilities in the neural network architecture.
@@ 2.3. Differentially Private Stochastic Gradient Descent (DPSGD)
Differential Privacy (DP)  (Dwork, 2006) establishes a mathematical framework that quantifies and bounds the privacy risk when operating on sensitive data. The framework ensures that statistical queries on a dataset remain nearly unchanged regardless of the inclusion or exclusion of any individual record, providing formal privacy guarantees. This privacy guarantee is parameterized by $\epsilon$ (privacy budget) and $\delta$ (failure probability), formalized through the following inequality:
@@
where $D$ and $D^{\prime}$ are datasets differing by one element, $M$ is a randomized algorithm, and $S$ is a subset of possible outputs. This inequality limits the probability of $M$ producing an output within $S$ when applied to $D$ is limited by the exponential of $\epsilon$. This probability is then multiplied by the probability of $M$ producing the same output from $D^{\prime}$. Additionally, it is adjusted by a small term $\delta$. In this context, $\epsilon$ controls the sensitivity of the output to variations in the input, where smaller values of $\epsilon$ indicate enhanced privacy by limiting the permissible changes in output probabilities. $\delta$, ideally a small value near zero, accounts for the rare cases where this privacy guarantee might fail, thus ensuring that changes to any single data point in $D$ minimally affect the output, reinforcing data privacy across the dataset.
DPSGD implements these theoretical guarantees in the context of neural network training by incorporating calibrated noise into the gradient computation process. In DPSGD, the key step involves perturbing the gradients computed during each training iteration with noise that is calibrated to the sensitivity of the function being optimized. This sensitivity measures how much the output of a function can change in response to changes in its input, which in the context of machine learning, translates to how much a single training example can influence the overall model parameters. The noise added is typically drawn from a Gaussian distribution  (Kassam, 2012), scaled according to $\epsilon$ and the desired level of privacy guarantee, $\delta$. The function of DPSGD can be expressed mathematically as: